The chief statistician of Aloha Company developed the following regression analysis:

Y = P49,272 + P1.78L + P2.68M

 

Where:

Y = total monthly manufacturing overhead cost

L = labor hours

M = machine hours

 

The measure of goodness of fit are good and no evidence of multicollinearity exists. Aloha Company will use 12,000 labor hours and 2,000 machine hours next month.

 

1.     Determine the total manufacturing overhead cost that Aloha Company should incur next month. 

2.     Aloha Company makes a product that has P6 in materials cost. It requires two hours of labor time and 30 minutes of machine time. Laborers earn P10 per hour. What is the product’s per unit variable manufacturing cost?

3.     Suppose that Aloha Company could reduce the labor time for the product described in requirement 2 by 30 minutes, to 1.50 hours. Machine time will remain the same. By how much would the per unit variable manufacturing cost fall?



1
Expert's answer
2021-11-08T17:29:27-0500

1.total manufacturing overhead cost


Y = P49,272 + P1.78L + P2.68M


Y=P49,272 + P1.78"\times" 12,000 + P2.68"\times" 2,000


Y=P75,992


2.Product’s per unit variable manufacturing cost

=P6"\times"2+P10"\times"0.5

=P17


3.By how much would the per unit variable manufacturing cost fall?

=P17-(P6"\times"1.5+P10"\times"0.5)

=P3